-2x+4y=10
-1x-4y=-13

Answer by DarkAgent13(1)

-2x+4y=10
-1x-4y=-13
4y=10+2x
-4y=-13+1x
y=5/2+1/2x
-1x-4(5/2+1/2x)=-13
-1x-10-2x=-13
-3x-10=-13
-3x=-3
x=1
-2(1)+4y=10
-2+4y=10
4y=12
y=3